# 2D-OOB: Attributing Data Contribution through Joint Valuation Framework

### v1

@@ 4.2 Cell fixation experiment
A naive strategy to handle cell-level outliers is to eliminate data points that contain outliers. This method, however, risks substantial data loss, particularly when outliers are scattered and data points are costly to collect. We instead consider a cell fixation experiment, where we assume that the ground-truth annotations of outlier cells can be restored with external expert knowledge. At each step, we “fix" a certain number of cells by substituting them with their ground-truth annotations, prioritizing cells that have the lowest valuations. Then we fit a logistic model and evaluate the model’s performance with a test set of $3000$ samples. It is important to note that correcting normal cells has no effect, whereas fixing outlier cells is expected to enhance the model’s performance. We adopt the same datasets and implementations as in Section 4.1.
Figure 3: Cell fixation experiment results (test accuracy curves) for 2D-OOB, 2D-KNN, and Random. We replace cells with their corresponding ground-truth annotations, starting with those cells assigned the lowest valuations. The results from $6$ datasets are presented, and additional results are provided in Appendix B.2. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB demonstrates a superior capability in accurately identifying and rectifying cell-level outliers.
@@ Results
Figure 4: Backdoor trigger detection rate curve for 2D-OOB, 2D-KNN, and Random. The panels A. and B. correspond to Trojan square and BadNets square, respectively. We prioritize cells within each poisoned sample, ranking from highest to lowest based on their valuations. The detection rate curve shows the average detection rate across all poisoned samples, and error bars represent a $95\%$ confidence interval based on 15 independent runs. 2D-OOB demonstrates superior performance in detecting the cells implanted with triggers.
@@ 4.3 Backdoor trigger detection
A common strategy of data poisoning attacks involves inserting a predefined trigger (e.g., a specific pixel pattern in an image) into a few training data [13, 5, 28]. These malicious manipulations can be challenging to detect as they only infect targeted samples. Even when poisoned data are present, it could be difficult to discern the cause of attacks since manually reviewing the images is expensive and time-consuming. In this experiment, we introduce a novel joint valuation task: detecting backdoor triggers in data poisoning attacks. Distinct from random outliers investigated previously, such cell contamination is targeted and deliberate.
Figure 5: Qualitative examples for 2D-OOB in the backdoor trigger detection task. Each pair of images shows a poisoned image and its cell valuation. The color of the heatmap indicates importance: red cells are more important than blue cells. The first two pairs consider the case the class “bird” is relabeled as “cat”, and the latter two pairs consider the case the class “deer” is relabeled as “cat”. The heatmaps clearly show that higher cell valuations predominantly concentrate on the regions containing triggers, while areas featuring actual objects receive lower valuations. This pattern suggests that 2D-OOB effectively captures the triggers as the impactful features responsible for the misclassification of the poisoned samples.
We consider two popular backdoor attack algorithms: BadNets [13] and Trojan Attack [28]. The poisoned samples, relabeled as the adversarial target class, are mixed up with the clean data in the training process. As a result, the model is trained to incorrectly treat the trigger as a main feature of the poisoned samples. At the test time, those inputs containing the trigger will be misclassified to the target class. In this context, our goal is to effectively pinpoint the triggers by recognizing them as influential features through our joint valuation framework.
Experimental setting We select $5$ pairs of CIFAR-10 classes. For each pair, we designate one as the target attack class and the other as the source class. The training dataset comprises $1000$ images. For each attack, we contaminate $15\%$ of the training samples from the source class and relabel them to the target class. Two types of attack triggers are implemented: Trojan square and BadNets square [13, 35, 28]. These triggers are placed in the lower right corner of the original images to minimize occlusion. Details of these attacks are available in the Appendix A.4. In our experiment, the ratio of poisoned cells is approximately $1\%$. We sample 25% features to build each weak learner.
@@ Results
We adopt the same detection scheme and baseline methods as in Section 4.1. Ideally, the poisoned cell should receive a high valuation based on the fact that such data point has been relabeled. We plot the detection rate curves of five datasets as shown in Figure 4. 2D-OOB significantly outperforms 2D-KNN in detecting both types of triggers. Overall, 2D-OOB achieves an average AUC of $0.95$ across all datasets and attack types, compared to $0.83$ for 2D-KNN.
@@
Figure 5 displays the heatmaps for poisoned samples based on cell valuations of 2D-OOB. Areas with higher cell valuations (marked as dark red color) precisely indicate the trigger location in these samples, illustrating the effectiveness of our detection. More examples are included in the Appendix B.3.
@@ 4.4 Ablation study
We conduct ablation studies on the cell-level outlier detection task, as outlined in Section 4.1, to examine the impact of the selection and number of weak learners on 2D-OOB estimations.
@@
Although our study primarily employs decision trees as weak learners, it is important to note that 2D-OOB is model-agnostic, enabling the use of any class of machine learning models as weak learners. We compare efficacy of decision trees, logistic regression, a single-layer MLP with 64 dimensions, and a two-layer MLP with 64 and 32 dimensions.
Table 2 presents a comparison of detection AUC across 12 datasets, indicating that 2D-OOB is not model-free. The selection of weak learners slightly affects the valuation results, with more complex models generally yielding better performance. Nonetheless, all variations of 2D-OOB outperform 2D-KNN, highlighting the significant advantages of the 2D-OOB approach.
Table 2: Ablation study results of weak learner types. The average and standard error of the AUC based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Results of 2D-KNN are added for comparison. Different types of weak learners lead to variations in the valuation results, with more complex models generally showing better performance.
@@
The number of weak learners
Increasing the number of weak learners allows for a greater number of data-feature subset pairs to be explored, potentially leading to more accurate estimates. However, we empirically observe that beyond a certain threshold, adding extra weak learners does not substantially enhance performance, indicating convergence of the estimation in Appendix B.4. As a summary, we vary the number of weak learners $B\in\{500,1000,3000\}$ and compare the cell-level outlier detection performance. Typically, when the number of weak learners is 1000, i.e., $B=1000$, it is sufficient to achieve converged estimates across different datasets.
Lastly, we present additional ablation study results for other key hyperparameters in Appendix B.4. Apart from the experiments discussed above, we showcase that marginalization of 2D-OOB can either match or surpass state-of-the-art data valuation methods on standard benchmarks in Appendix D.
@@ 6 Conclusion
We propose 2D-OOB, an efficient joint valuation framework that assigns a score to each cell in a dataset, thereby facilitating a finer attribution of data contribution and enabling a deeper understanding of datasets. Through comprehensive experiments, we show that 2D-OOB is computationally efficient and competitive over state-of-the-art methods in both joint valuation tasks.
@@
In the supplementary materials, we provide implementation details, additional experimental results, rigorous formalized proofs and data valuation experiment results. Code repository can be found at https://github.com/YifanSun99/2d-oob.
@@ A.1 Datasets
We use $12$ binary classification datasets obtained from OpenML [11]. A summary of all the datasets is provided in Table 3. These datasets are used in Section 4.14.2, and Appendix D.
For each dataset, we first employ a standard normalization procedure, where each feature is normalized to have zero mean and unit standard deviation. After preprocessing, we randomly partition a subset of the data into two non-overlapping sets: a training dataset and a test dataset, which consists of $1000$ and $3000$ samples respectively. The training dataset is used to obtain the joint (or marginal) valuation for each cell (or data point). The test dataset is exclusively used for cell fixation (or point removal) experiments when evaluating the test accuracy. Note that for methods that need a validation dataset such as KNNShapley and DataShapley, we additionally sample a separate validation dataset (disjoint from training dataset and test dataset) to evaluate the utility function. The size of the validation dataset is set to $10\%$ of the training sample size.
@@
We create datasets by pairing CIFAR-10 classes, each pair consisting of a target attack class and a source class. The training and test dataset comprises $1000$ and $2000$ samples respectively. The size of the validation dataset is set to $10\%$ of the training sample size. To manage the computational challenges posed by the baseline method, we employ the super-pixel technique to transform the (32,32,3) image into a $256$-dimensional vector. Specifically, we first average the pixel values across three channels for each pixel. Then, we partition these transformed images into equally sized $2\times 2$ grids. In each grid, we use average pooling to reduce the pixel values to a single cell value. These cell values are then arranged into a flattened input vector. We annotate a cell as poisoned if at least $25\%$ of its corresponding grid area contains the trigger.
@@ A.2 Implementation details for different methods
2D-OOB involves fitting a subset random forest model with $B=1000$ decision trees based on the package “scikit-learn”. When constructing each decision tree, we fix the feature subset size ratio as $0.5$. Ablations on the hyperparameters can be found in Appendix B.4. For Section 4.3 and Appendix D, we simply adopt $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))$. For Section 4.1 and 4.2, we further calculate the normalized negative L2 distance between covariates and the class-specific mean in the bootstrap dataset, denoted as $d_{norm}$. Then we use $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))+d_{norm}$.
@@
2D-KNN employs KNN as a surrogate model to approximate 2D-Shapley. We set the number of nearest neighbors as $10$ and the number of permutations as $1000$. The hyperparameters are selected based on convergence behavior and we determine the run time until the values converge.
@@ A.3 Implementation details for cell-level outlier generation
Following Du et al. [8] and Liu et al. [29], we replace a given cell with the outlier value. Here, the outlier value is randomly generated from the two-sided “tails” of the Gaussian distribution with the column mean and standard deviation, where the probability of the two-sided tail area is set to be $1$%. $4\%$ ($20\%\times 20\%$) of the cells in total are replaced with the corresponding outlier value.
@@ A.4 Implementation details for backdoor trigger generation
Following the prior work [13, 28], we generate the BadNets square and the Trojan square trigger. For BadNets, we adopt the implementation in Nicolae et al. [34]. For Trojan Attack, we use a pretrained ResNet18 model on CIFAR-10 dataset and employ the implementation in Pang et al. [35]. For each attack, we evaluate its effectiveness by training a decision tree model on the poisoned dataset. The accuracy on a clean test set remains nearly unchanged compared to the model trained on an uncontaminated training set, while the attack success rate on a hold-out poisoned test sample set is guaranteed to exceed $75\%$.

### v2

@@ 4.2 Cell fixation experiment
A naive strategy to handle cell-level outliers is to eliminate data points that contain outliers. This method, however, risks substantial data loss, particularly when outliers are scattered and data points are costly to collect. We instead consider a cell fixation experiment, where we assume that the ground-truth annotations of outlier cells can be restored with external expert knowledge. At each step, we “fix" a certain number of cells by substituting them with their ground-truth annotations, prioritizing cells that have the lowest valuations. Then we fit a logistic model [^1] Logistic regression is chosen because it is a simple yet powerful machine learning model commonly used to test data separability, which is a standard practice in this field [16, 30]. and evaluate the model’s performance with a test set of $3000$ samples. It is important to note that correcting normal cells has no effect, whereas fixing outlier cells is expected to enhance the model’s performance. We adopt the same datasets and implementations as in Section 4.1.
Figure 3: Cell fixation experiment results (test accuracy curves) for 2D-OOB, 2D-KNN, and Random. We replace cells with their ground-truth annotations, starting with those cells assigned the lowest valuations. The results for $6$ datasets are presented, and additional results for other datasets are available in Appendix B.2. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB demonstrates a superior capability in accurately identifying and rectifying cell-level outliers.
@@ Results
Figure 4: Backdoor trigger detection rate curves for 2D-OOB, 2D-KNN, and Random. Panels A (top) and B (bottom) correspond to the Trojan square and BadNets square, respectively. We inspect the cells within each poisoned sample in descending order of their valuation scores. The detection rate curve shows the average detection rate across all poisoned samples, with error bars representing a $95\%$ confidence interval based on $15$ independent runs. 2D-OOB demonstrates superior performance in detecting the cells implanted with triggers.
@@ 4.3 Backdoor trigger detection
A common strategy of data poisoning attacks involves inserting a predefined trigger (e.g., a specific pixel pattern in an image) into a subset of the training data [13, 5, 29]. These malicious manipulations can be challenging to detect as they only infect specific, targeted samples. Even when poisoned data are present, it could be difficult to discern the root cause of the attacks, since manually reviewing the images is expensive and time-consuming. In this experiment, we introduce a novel task enabled by the joint valuation framework: localizing backdoor triggers in data poisoning attacks. Distinct from the random outliers investigated previously, this type of cell contamination is targeted and deliberate.
Figure 5: Qualitative examples for 2D-OOB in the backdoor trigger detection task. Each pair of images consists of a poisoned image and its corresponding cell valuation heatmap. The color of the heatmap indicates importance, with red cells representing higher importance and blue cells representing lower importance. In the first two pairs, the class ‘bird” is relabeled as “cat”, while in the latter two pairs, the class “deer” is relabeled as “cat”. The heatmaps clearly show that higher cell valuations are predominantly concentrated in the regions containing triggers, while areas featuring actual objects receive lower valuations. This pattern suggests that 2D-OOB effectively captures the triggers as the impactful features responsible for the misclassification of the poisoned samples.
We consider two popular backdoor attack algorithms: BadNets [13] and Trojan Attack [29]. During training, the poisoned samples, relabeled as the adversarial target class, are mixed up with clean data. As a result, the model learns to incorrectly associate the trigger with the target class. At test time, inputs containing the trigger are misclassified to the target class. In this context, our goal is to effectively pinpoint the triggers by recognizing them as influential features through our joint valuation framework.
Experimental setting We select 5 pairs of classes from CIFAR-10 [22]. For each pair, one class is designated as the target attack class, while the other serves as the source class. The training dataset comprises $1000$ images. For each attack, we contaminate $15\%$ of the training samples from the source class and relabel them to the target class. Two types of attack triggers are implemented: the Trojan square and the BadNets square [13, 36, 29]. These triggers are placed in the lower right corner of the original images to minimize occlusion. Further details about these attacks are available in Appendix A.4. In our experiment, the ratio of poisoned cells is approximately $1\%$, and each weak learner in the subset bagging model is built by sampling $25$% of the features.
@@ Results
We adopt the same detection scheme and baseline methods as in Section 4.1. Ideally, the poisoned cells should receive high valuation scores given that such data points have been relabeled. We present the detection rate curves for five datasets in Figure 4. 2D-OOB significantly outperforms 2D-KNN in detecting both types of triggers. Overall, 2D-OOB achieves an average detection AUC of $0.95$ across all datasets and attack types, compared to $0.83$ for 2D-KNN. It is worth noting that conventional data valuation methods are fundamentally unable to localize backdoor triggers; at most, they can only identify poisoned data points.
@@
Figure 5 displays the heatmaps for poisoned samples based on cell valuations of 2D-OOB. Areas with higher cell valuations (marked in dark red) precisely indicate the trigger locations within these samples, demonstrating the effectiveness of our joint valuation framework. Additional examples can be found in Appendix B.3.
@@ 4.4 Ablation study
We conduct ablation studies on the cell-level outlier detection task, as outlined in Section 4.1, to examine the impact of key hyperparameters on 2D-OOB estimations, including the selection and number of weak learners, as well as the feature subset ratio.
@@
Although our study primarily employs decision trees as weak learners, it is important to note that 2D-OOB is model-agnostic, enabling the use of any class of machine learning models as weak learners. Specifically, we examine decision trees, logistic regression, a single-layer MLP with $64$ units, and a two-layer MLP with $64$ and $32$ units, respectively, as weak learners to compare their performance.
Table 2 presents a comparison of detection AUC across $12$ datasets with different choices of weak learners, indicating that 2D-OOB is not model-free. The selection of weak learners slightly affects the valuation results, with more complex models generally yielding better performance. Nonetheless, all variations of 2D-OOB outperform 2D-KNN, highlighting the clear advantages of the 2D-OOB approach.
Table 2: Ablation study results of weak learner types. The average and standard error of the detection AUC based on $30$ independent experiments are denoted by “average $\pm$ standard error”. Results from 2D-KNN are included for comparison. The choice of weak learner leads to variations in cell values, yet the performance of the detection task remains robust.
@@
Number of weak learners
Increasing the number of weak learners allows for a greater number of data-feature subset pairs to be explored, potentially leading to more accurate estimates. However, as shown in Table 3(a), when we increase the number of base learners from $500$ to $3000$, the detection AUC for each dataset remains relatively unchanged, suggesting convergence of the estimation beyond a certain threshold. Typically, $1000$ base learners are sufficient to achieve an equitable joint valuation.
Feature subset ratio $K/d$
The feature subset ratio $K/d$ refers to the fraction of the total number of features $d$ that are randomly selected to build each weak learner, where $K$ is the number of selected features. In previous experiments, we used a fixed ratio of $0.50$ (unless otherwise specified). To further investigate the impact of this ratio, we now test two additional values: $0.25$ and $0.75$. The results in Table 3(b) suggest that in general, the joint valuation capacity of our method is robust to the choice of feature subset ratio.
Table 3: Ablation study results of (a) the number of base learners $B$ and (b) the feature subset ratio $K/d$. The average and standard error of the detection AUC based on $30$ independent runs are denoted by “average $\pm$ standard error.” (a) Increasing the number of base learners from $1000$ to $3000$ does not yield a notable performance improvement. (b) Our method’s joint valuation capacity remains relatively stable regardless of the selected feature subset ratio.
(a) Ablation on the number of base learners $B$.
Dataset AUC $\uparrow$
$B=500$ $B=1000$ $B=3000$
lawschool 0.86 $\pm$ 0.0035 0.88 $\pm$ 0.0027 0.88 $\pm$ 0.0026
electricity 0.77 $\pm$ 0.0062 0.77 $\pm$ 0.0072 0.77 $\pm$ 0.0070
fried 0.87 $\pm$ 0.0022 0.91 $\pm$ 0.0015 0.91 $\pm$ 0.0014
2dplanes 0.87 $\pm$ 0.0016 0.87 $\pm$ 0.0015 0.87 $\pm$ 0.0015
creditcard 0.72 $\pm$ 0.0025 0.72 $\pm$ 0.0028 0.72 $\pm$ 0.0028
pol 0.78 $\pm$ 0.0022 0.82 $\pm$ 0.0014 0.82 $\pm$ 0.0014
MiniBooNE 0.77 $\pm$ 0.0042 0.77 $\pm$ 0.0058 0.77 $\pm$ 0.0058
jannis 0.78 $\pm$ 0.0045 0.83 $\pm$ 0.0042 0.83 $\pm$ 0.0039
nomao 0.79 $\pm$ 0.0018 0.79 $\pm$ 0.0021 0.79 $\pm$ 0.0020
vehicle_sensIT 0.80 $\pm$ 0.0021 0.81 $\pm$ 0.0014 0.81 $\pm$ 0.0014
gas_drift 0.86 $\pm$ 0.0007 0.86 $\pm$ 0.0010 0.86 $\pm$ 0.0010
musk 0.88 $\pm$ 0.0008 0.88 $\pm$ 0.0008 0.88 $\pm$ 0.0008
Average 0.81 0.83 0.83
(b) Ablation on feature subset ratio $K/d$.
Dataset AUC $\uparrow$
$K/d=0.25$ $K/d=0.50$ $K/d=0.75$
lawschool 0.86$\pm$ 0.0026 0.88$\pm$ 0.0027 0.88$\pm$ 0.0024
electricity 0.79$\pm$ 0.0070 0.77$\pm$ 0.0072 0.73$\pm$ 0.0070
fried 0.86$\pm$ 0.0024 0.91$\pm$ 0.0015 0.89$\pm$ 0.0007
2dplanes 0.82$\pm$ 0.0015 0.87$\pm$ 0.0015 0.88$\pm$ 0.0014
creditcard 0.73$\pm$ 0.0029 0.72$\pm$ 0.0028 0.71$\pm$ 0.0028
pol 0.66$\pm$ 0.0031 0.82$\pm$ 0.0014 0.82$\pm$ 0.0014
MiniBooNE 0.78$\pm$ 0.0076 0.77$\pm$ 0.0058 0.77$\pm$ 0.0049
jannis 0.84$\pm$ 0.0035 0.83$\pm$ 0.0042 0.82$\pm$ 0.0043
nomao 0.79$\pm$ 0.0019 0.79$\pm$ 0.0021 0.78$\pm$ 0.0021
vehicle_sensIT 0.81$\pm$ 0.0014 0.81$\pm$ 0.0014 0.80$\pm$ 0.0015
gas_drift 0.88$\pm$ 0.0009 0.86$\pm$ 0.0010 0.86$\pm$ 0.0009
musk 0.89$\pm$ 0.0008 0.88$\pm$ 0.0008 0.88$\pm$ 0.0008
Average 0.81 0.83 0.82
Apart from the experiments discussed above, we showcase that marginalization of 2D-OOB can either match or surpass state-of-the-art data valuation methods on standard benchmarks in Appendix D.
@@ 6 Conclusion
We propose 2D-OOB, an efficient joint valuation framework that assigns a score to each cell in a dataset, thereby facilitating finer attribution of data contributions and enabling a deeper understanding of the dataset. Through comprehensive experiments, we show that 2D-OOB is computationally efficient and competitive over state-of-the-art methods in multiple joint valuation use cases.
Discussion
We emphasize that the primary objective of the joint valuation framework is to evaluate the quality of cells within the dataset, rather than to optimize model performance. The model used serves as a proxy for this evaluation, and it is important to note that a high-performing machine learning method does not necessarily ensure a justified valuation framework.
@@ Discussion
Acknowledgement
We acknowledge computing resources from Columbia University’s Shared Research Computing Facility project, which is supported by NIH Research Facility Improvement Grant 1G20RR030893-01, and associated funds from the New York State Empire State Development, Division of Science Technology and Innovation (NYSTAR) Contract C090171, both awarded April 15, 2010.
@@ Acknowledgement
In the supplementary materials, we provide implementation details, additional experimental results, rigorous formalized proofs and data valuation experiment results. Code repository can be found at https://github.com/yifansun99/2D-OOB-Joint-Valuation.
@@ A.1 Datasets
We use $12$ binary classification datasets obtained from OpenML [11]. A summary of all the datasets is provided in Table 4. These datasets are used in Section 4.14.24.4, and Appendix D.
For each dataset, we first employ a standard normalization procedure, where each feature is normalized to have zero mean and unit standard deviation. After preprocessing, we randomly partition a subset of the data into two non-overlapping sets: a training dataset and a test dataset, which consists of $1000$ and $3000$ samples respectively. The training dataset is used to obtain the joint (or marginal) valuation for each cell (or data point). The test dataset is exclusively used for the cell fixation (or point removal) experiment when evaluating the test accuracy. Note that for methods that need a validation dataset such as KNNShapley and DataShapley, we additionally sample a separate validation dataset (disjoint from training dataset and test dataset) to evaluate the utility function. The size of the validation dataset is set to $10\%$ of the training sample size.
@@
We create datasets by pairing CIFAR-10 classes, each pair consisting of a target attack class and a source class. The training and test dataset comprises $1000$ and $2000$ samples, respectively. The size of the validation dataset is set to $10\%$ of the training sample size. To manage the computational challenges posed by the baseline method, we employ the super-pixel technique to transform the ($32$,$32$,$3$) image into a $256$-dimensional vector. Specifically, we first average the pixel values across the three channels for each pixel. Then, we partition these transformed images into equal-sized $2\times 2$ grids. Average pooling is applied within each grid to reduce pixel values to a single cell value, which is then arranged into a flattened input vector. A cell is annotated as poisoned if at least $25\%$ of its corresponding grid area contains the trigger.
@@ A.2 Implementation details for different methods
2D-OOB involves fitting a subset random forest model with $B=1000$ decision trees based on the package “scikit-learn”. When constructing each decision tree, we fix the feature subset size ratio as $0.5$ (unless otherwise specified). For Section 4.3 and Appendix D, we simply adopt $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))$. For Section 4.1 and 4.2, we further calculate the normalized negative $L_{2}$ distance between covariates and the class-specific mean in the bootstrap dataset, denoted as $d_{norm}(x_{i},y_{i})$. Then we use $T(y_{i},\hat{f}(x_{i,S_{b}}))=\mathds{1}(y_{i}=\hat{f}(x_{i,S_{b}}))+d_{norm}(x_{i},y_{i})$.
@@
2D-KNN employs KNN as a surrogate model to approximate 2D-Shapley. We set the number of nearest neighbors as $10$ and the number of permutations as $1000$. The hyperparameters are selected based on convergence behavior, and the run time is measured until the values converge.
@@ A.3 Implementation details for cell-level outlier generation
Following Du et al. [8] and Liu et al. [30], we replace a given cell with an outlier value. Here, the outlier value is randomly generated from the two-sided “tails” of the Gaussian distribution fitted to the column’s mean and standard deviation, where the probability of the two-sided tail area is set to be $1$%. In total, $4\%$ ($20\%\times 20\%$) of the cells are replaced with the corresponding outlier values.
@@ A.4 Implementation details for backdoor trigger generation
Following the prior work [13, 29], we generate the BadNets square and the Trojan square trigger. For BadNets, we adopt the implementation in Nicolae et al. [35]. For Trojan Attack, we use a pretrained ResNet-18 model on the CIFAR-10 dataset and employ the implementation in Pang et al. [36]. For each attack, we evaluate its effectiveness by training a decision tree model on the poisoned dataset. The accuracy on a clean test set remains nearly unchanged compared to a model trained on an uncontaminated training set, while the attack success rate on a held-out poisoned test set is guaranteed to exceed $75\%$.